# London WC1E 6BT, United Kingdom

### v1

@@ 1 Introduction
Heavy Neutral Leptons (HNLs) are amongst the most popular exotic particles thought to exist beyond the particle content of the Standard Model (SM). Assumed to be fermions that are neutral under the SM gauge symmetries, they act as heavy right-handed neutrinos and, if they are of Majorana nature, trigger a seesaw mechanism of light neutrino mass generation Minkowski:1977sc ; Mohapatra:1979ia ; Gell-Mann:1979vob ; Yanagida:1979as ; Schechter:1980gr . In this context, they may also help to understand the matter-antimatter asymmetry of the universe, as their $CP$-violating decays or oscillations can generate a baryon asymmetry through leptogenesis Fukugita:1986hr . HNLs may thus help to explain why the active neutrinos are so much lighter than the other SM fermions, but not massless as required from observing oscillations ParticleDataGroup:2024cfk ; Esteban:2020cvm , and the origin of matter in the universe.
@@
In the minimal scenario, an HNL $N$ couples to the SM only through its mixing $U_{\alpha N}$ ($\alpha=e,\mu,\tau$) with the active neutrinos, induced by the Yukawa coupling of the HNL with the SM lepton doublets $L_{\alpha}$ and Higgs doublet $H$. The HNL then also participates in charged and neutral lepton currents. There are numerous past, current and planned future searches for HNLs based on this, over a wide range of HNL masses $m_{N}$; from the eV-scale where HNLs can be tested in oscillations, over keV and MeV scales mainly probed in nuclear processes such as $\beta$ decay, MeV to GeV in beam dump and meson decays, to electroweak scale masses and above probed in colliders. A recent overview of current and future searches is provided in Bolton:2019pcu ; Abdullahi:2022jlv and a global study of current constraints on HNLs can be found in Chrzaszcz:2019inj . The mass range $100~{}\text{MeV}\lesssim m_{N}\lesssim 100$ GeV has especially been targeted in searches for HNLs, since in this mass range HNLs are naturally long-lived for the small active-sterile mixing strengths expected for successful neutrino mass generation. This results in macroscopic decay lengths and thus displaced decay vertices that can be looked for with a high sensitivity.
@@
Being long-lived, HNLs also affect the early history of the universe beyond their role in leptogenesis. If HNLs had been in thermal equilibrium, and they decay around or later than about a second after the big bang, the produced particles affect Big Bang Nucleosynthesis (BBN) Boyarsky:2021yoh ; GAMBITCosmologyWorkgroup:2020htv . In the absence of other channels, the primary decay of HNLs to SM particles is hadronically Atre:2009rg ; Coloma:2020lgy . The mesons produced through these decays interact strongly with the protons and neutrons of the primordial plasma, thereby modifying the $p\leftrightarrow n$ conversion ratio which sets the primordial abundances. Most notably, the primordial abundance of 4He, precisely measured as $Y_{p}=0.2436\pm 0.0038$ Hsyu:2020uqb , would be mostly affected since most primordial neutrons and protons end up forming 4He. Such considerations disfavour HNLs with masses $m_{N}\lesssim 1$ GeV, for small active-sterile mixing strengths required by current constraints and expected for light neutrino mass generation. Even lighter and longer-lived $N$ are likewise disfavoured as they will act as additional degrees of freedom, or inject them through their decays, and may overclose the universe.
@@
The above considerations apply for the minimal scenario where the HNL only couples via the active-sterile mixing. Instead, HNLs may also interact with an expanded exotic sector within specific beyond-the-SM scenarios. Often, these are high-scale scenarios where the HNL is charged under an additional gauge force, such as $B-L$ Davidson:1978pm ; Marshak:1979fm ; Mohapatra:1980qe ; Davidson:1987mh ; Buchmuller:1991ce and left-right-symmetric Pati:1973rp ; Pati:1973uk ; Pati:1974yy ; Mohapatra:1974gc ; Senjanovic:1975rk ; Mohapatra:1980qe models. Less explored is the possibility of coupling the HNL to a lighter exotic sector. We here explore such a scenario where the HNL couples to a light, exotic pseudoscalar $a$, akin to an axion-like particle (ALP) Marsh:2015xka ; Chadha-Day:2021szb ; Marsh:2023tep . We will refer to this dark scalar as ALP in the following, and it is dark in the sense that it couples only to the HNL in the first instance, with interactions to the SM suppressed by the active-sterile mixing and through loops.
@@ 2.1 Lagrangian
$\mathcal{M}_{\nu}=\begin{pmatrix}0&\mathcal{M}_{D}\\ \mathcal{M}_{D}^{T}&\mathcal{M}_{R}\end{pmatrix},$ (2)
@@
$U=\begin{pmatrix}U_{\nu\nu}&U_{\nu N}\\ U_{N\nu}&U_{NN}\end{pmatrix},$ (3)
@@
$U^{\dagger}\mathcal{M}_{\nu}U^{\ast}=U^{\dagger}\cdot\begin{pmatrix}0&\mathcal{M}_{D}\\ \mathcal{M}_{D}^{T}&\mathcal{M}_{R}\end{pmatrix}\cdot U^{\ast}=\begin{pmatrix}(m_{\nu})_{3\times 3}&0\\ 0&(M_{N})_{\mathcal{N}\times\mathcal{N}}\end{pmatrix}.$ (4)
@@
In addition to the regular seesaw sector, we also assume that the ALPs $a$ couples with the HNL $N$, but not with other SM particles directly. Hence, due to the pseudoscalar nature of ALPs, the Lagrangian for the ALP-HNL interaction is given by the derivative coupling Gola:2021abm in the mass basis as
$\mathcal{L}_{aNN}=\sum_{\kappa=1}^{\mathcal{N}}\frac{1}{f_{a}}(\partial_{\mu}a)\bar{N}_{\kappa}\gamma^{\mu}\gamma_{5}N_{\kappa}=-\sum_{\kappa=1}^{\mathcal{N}}\frac{2i}{f_{a}}m_{N_{\kappa}}a\bar{N}_{\kappa}\gamma_{5}N_{\kappa},$ (6)
where $f_{a}$ is the ALP decay constant and $m_{N_{\kappa}}$ are the eigenvalues of $M_{N}$. The second equality holds by applying the equation of motion for the HNLs and removing a total derivative after integration by parts Gelmini:1982zz . Due to the small active-sterile mixing $U_{\nu N}=U_{N\nu}\equiv\sqrt{m_{\nu}M_{N}^{-1}}$, an interaction between the ALP and the active neutrinos $\nu_{\lambda}$ is induced. In the mass basis of the active neutrinos and HNLs, we can write the interaction Lagrangian between the ALP and active neutrinos as
@@
$\approx-\sum_{\lambda,\lambda^{\prime}=1}^{3}\frac{2i}{f_{a}}m_{N}a|U_{N\nu}U_{\nu\nu}^{\ast}|^{2}_{\lambda\lambda^{\prime}}\bar{\nu}_{\lambda}\gamma_{5}\nu_{\lambda^{\prime}}$ (7)
@@
$\mathcal{L}_{aN\nu}=-\sum_{\lambda^{\prime}=1}^{3}\sum_{\kappa=1}^{\mathcal{N}}\frac{2i}{f_{a}}m_{N_{\kappa}}a(U_{N\nu}^{\ast}U_{\nu\nu})_{\kappa\lambda^{\prime}}\bar{N}_{\kappa}\gamma_{5}\nu_{\lambda^{\prime}}.$ (8)
While at least two HNLs are needed to explain all active neutrino masses and mixing, we are mainly interested in elucidating the principle effects in our framework. For simplicity, by considering a single HNL, $N_{1}\equiv N$, we can acquire all the important information from the framework using this simplified viewpoint. Furthermore, we consider the active-sterile mixing of $N_{1}$ with a single active neutrino $\nu_{1}$ which is mostly electron-type neutrino i.e., $\nu_{1}\equiv\nu_{e}$, without loss of generality. From now on, we take $|U_{\nu\nu}|^{2}\sim\mathcal{O}(1)\gg|U_{\nu N}|^{2}=|U_{N\nu}|^{2}$ and $U_{\nu N}=U_{eN}$. Equations (6) and (7) then take the form
@@
$\mathcal{L}_{a\nu\nu}$ $=-\frac{2i}{f_{a}}m_{N}|U_{eN}|^{2}a\bar{\nu}_{e}\gamma_{5}\nu_{e}=-\frac{2i}{f_{a}}m_{\nu}a\bar{\nu}_{e}\gamma_{5}\nu_{e},$ (9)
@@ 2.2 HNL decays
A Majorana HNL with a mass of a few hundred MeV to GeV decays via various channels into SM particles. All these decays are four-fermion interactions mediated by either a $Z$ or $W$ boson. We will briefly discuss these SM decay channels of HNLs here with a detailed discussion found in Atre:2009rg ; Coloma:2020lgy . In our framework with only one HNL species that couples with the first generation only, we have the corresponding decay channels of HNLs as:
@@
$\Gamma^{\nu_{e}\ell^{-}\ell^{+}}=|U_{eN}|^{2}\frac{G_{F}^{2}m_{N}^{5}}{96\pi^{3}}\left[\left(C_{1}+2\sin^{2}\theta_{W}\delta_{e,\ell}\right)f_{1}(x_{\ell})+\left(C_{2}+\sin^{2}\theta_{W}\delta_{e,\ell}\right)f_{2}(x_{\ell})\right],$ (10)
@@
$C_{1}=\frac{1}{4}(1-4\sin^{2}\theta_{W}+8\sin^{4}\theta_{W}),\qquad C_{2}=\frac{1}{2}(-\sin^{2}\theta_{W}+2\sin^{4}\theta_{W}),$ (11)
@@
$f_{2}(x)$ $=4\left[x^{2}(2+10x^{2}-12x^{4})\sqrt{1-4x^{2}}+6x^{4}(1-2x^{2}+2x^{4})L(x)\right],$ (12)
@@
$L(x)=\ln\left[\frac{1-3x^{2}-(1-x^{2})\sqrt{1-4x^{2}}}{x^{2}(1+\sqrt{1-4x^{2}})}\right].$ (13)
@@
$\Gamma^{e^{-}\ell^{+}\nu_{\ell}}=|U_{eN}|^{2}\frac{G_{F}^{2}m_{N}^{5}}{192\pi^{3}}\left[1-8x_{\ell}^{2}+8x_{\ell}^{6}-x_{\ell}^{8}-12x_{\ell}^{4}\ln(x_{\ell}^{2})\right].$ (14)
@@
$\Gamma^{P\nu_{e}}=\frac{G_{F}^{2}m_{N}^{3}}{32\pi}f_{P}^{2}|U_{eN}|^{2}(1-x_{P}^{2})^{2},$ (16)
with the meson decay constant $f_{P}$ whose values we have taken from Coloma:2020lgy and $x_{P}=m_{P}/m_{N}$.
@@
$\Gamma^{P^{+}e^{-}}=\frac{G_{F}^{2}m_{N}^{3}}{16\pi}f_{P}^{2}|U_{eN}|^{2}|V_{qq^{\prime}}|^{2}\lambda^{1/2}(1,x_{P}^{2},x_{e}^{2})\left[1-x_{P}^{2}-x_{e}^{2}(2+x_{P}^{2}-x_{e}^{2})\right].$ (17)
@@
$\Gamma^{V\nu_{e}}=\frac{G_{F}^{2}m_{N}^{3}}{32\pi m_{V}^{2}}f_{V}^{2}\kappa_{V}^{2}|U_{eN}|^{2}(1+2x_{V}^{2})(1-x_{V}^{2})^{2}.$ (18)
@@
$\Gamma^{V^{+}e^{-}}=\frac{G_{F}^{2}m_{N}^{3}}{16\pi m_{V^{\pm}}^{2}}f_{V}^{2}|U_{eN}|^{2}|V_{qq^{\prime}}|^{2}\lambda^{1/2}(1,x_{V}^{2},x_{e}^{2})\left[(1-x_{V}^{2})(1+2x_{V}^{2})+x_{e}^{2}(x_{V}^{2}+x_{e}^{2}-2)\right].$ (19)
@@
$+\sum_{P}\Gamma^{P\nu_{e}}+\sum_{P}2\Gamma^{P^{+}e^{-}}+\sum_{V}\Gamma^{V\nu_{e}}+\sum_{V}2\Gamma^{V^{+}e^{-}}.$ (20)
@@
Due to the interaction in Eq. (9), a new decay channel for HNLs is also possible, where the HNL decays as $N\to a\nu$ via the active-sterile mixing at tree level. The decay width, in the limit where $m_{\nu}\ll m_{N}$, can be written as
$\Gamma^{N\to a\nu}=\frac{|U_{eN}|^{2}m_{N}^{3}}{4\pi f_{a}^{2}}\sqrt{1+\left(\frac{m_{a}}{m_{N}}\right)^{2}}\left[1-\left(\frac{m_{a}}{m_{N}}\right)^{2}\right]^{3/2}\approx\frac{|U_{eN}|^{2}m_{N}^{3}}{4\pi f_{a}^{2}},$ (21)
@@
$\tau_{N\to a\nu}\approx 8.6\times 10^{-4}~{}\text{s}\times\left(\frac{f_{a}}{1~{}\text{TeV}}\right)^{2}\times\left(\frac{10^{-14}}{|U_{eN}|^{2}}\right)\times\left(\frac{1~{}\text{GeV}}{m_{N}}\right)^{3}.$ (22)
The total HNL decay width to SM particles can be approximately written as Bolton:2019pcu
@@
$\approx\left(\frac{6f_{M}^{2}}{\pi}+\frac{m_{N}^{2}}{20\pi^{3}}\right)m_{N}^{3}|U_{eN}|^{2}G_{F}^{2},$ (23)
where $f_{M}\approx\mathcal{O}(0.1)$ GeV corresponds to the typical decay constant of the produced pseudoscalar or vector meson. Comparing this with Eq. (21), it is clear that the $N\to a\nu$ channel will be dominant if
$m_{N}<\frac{\sqrt{5|1-24G_{F}^{2}f_{M}^{2}f_{a}^{2}|}\pi}{f_{a}G_{F}},$ (24)
@@
The branching ratios of the HNL decaying to various particles as a function of mass of the HNL for two different values of $f_{a}$ are shown in Fig. 1. In this work, as we are mostly interested in the regime where $m_{a}\ll m_{N}$, correspondingly we have chosen $m_{a}=1$ keV and varied $m_{N}\in[10^{-3}-50]$ GeV range. In the left panel, we have considered $f_{a}=1$ TeV where the Br$(N\rightarrow a\nu)\approx 100\%$ (solid green) throughout the entire mass range of HNL, while the behaviour of different SM decay channels are shown as, $N\rightarrow\nu\nu\nu$ (solid black), $N\rightarrow\nu_{e}(e^{+}e^{-}+\mu^{+}\mu^{-}+\tau^{+}\tau^{-})$ (solid orange), $N\rightarrow e^{-}(\nu_{\mu}\mu^{+}+\nu_{\tau}\tau^{+})$ (solid magenta), different pseudoscalar ($P$) and vector ($V$) mesonic channels (which become relevant for $m_{N}\geq m_{\pi}$) as $N\rightarrow P\nu$ (solid red), $N\rightarrow Pe$ (dashed brown), $N\rightarrow V\nu$ (solid cyan), $N\rightarrow Ve$ (dashed blue). In the right panel of the figure we have taken $f_{a}=10^{2.5}$ TeV, for which the coupling of the ALP to HNL significantly decreases. Here, for lower mass of HNL the axionic decay channel still dominates the scenario, as expected from eq. (24), while for $m_{N}>m_{\pi}$, as the mesonic channels come into picture, the branching ratio to axionic channel drops significantly as compared to the SM decay channels.

### v2

# Relaxing Limits from Big Bang Nucleosynthesis on Heavy Neutral Leptons with Axion-like Particles

@@ 1 Introduction
Heavy Neutral Leptons (HNLs) are amongst the most popular exotic particles thought to exist beyond the particle content of the Standard Model (SM). Assumed to be fermions that are neutral under the SM gauge symmetries, they act as heavy right-handed neutrinos and, if they are of Majorana nature, trigger a seesaw mechanism of light neutrino mass generation [1, 2, 3, 4, 5]. In this context, they may also help to understand the matter-antimatter asymmetry of the universe, as their $CP$-violating decays or oscillations can generate a baryon asymmetry through leptogenesis [6]. HNLs may thus help to explain why the active neutrinos are so much lighter than the other SM fermions, but not massless as required from observing oscillations [7, 8], and the origin of matter in the universe.
@@
In the minimal scenario, an HNL $N$ couples to the SM only through its mixing $U_{\alpha N}$ ($\alpha=e,\mu,\tau$) with the active neutrinos, induced by the Yukawa coupling of the HNL with the SM lepton doublets $L_{\alpha}$ and Higgs doublet $H$. The HNL then also participates in charged and neutral lepton currents. There are numerous past, current and planned future searches for HNLs based on this, over a wide range of HNL masses $m_{N}$; from the eV-scale where HNLs can be tested in oscillations, over keV and MeV scales mainly probed in nuclear processes such as $\beta$ decay, MeV to GeV in beam dump and meson decays, to electroweak scale masses and above probed in colliders. A recent overview of current and future searches is provided in [9, 10] and a global study of current constraints on HNLs can be found in [11]. The mass range $100~{}\text{MeV}\lesssim m_{N}\lesssim 100$ GeV has especially been targeted in searches for HNLs, since in this mass range HNLs are naturally long-lived for the small active-sterile mixing strengths expected for successful neutrino mass generation. This results in macroscopic decay lengths and thus displaced decay vertices that can be looked for with a high sensitivity.
@@
Being long-lived, HNLs also affect the early history of the universe beyond their role in leptogenesis. If HNLs had been in thermal equilibrium, and they decay around or later than about a second after the big bang, the produced particles affect Big Bang Nucleosynthesis (BBN) [12, 13]. In the absence of other channels, the primary decay of HNLs to SM particles is hadronically [14, 15]. The mesons produced through these decays interact strongly with the protons and neutrons of the primordial plasma, thereby modifying the $p\leftrightarrow n$ conversion ratio which sets the primordial abundances. Most notably, the primordial abundance of 4He, precisely measured as $Y_{p}=0.2436\pm 0.0038$ [16], would be mostly affected since most primordial neutrons and protons end up forming 4He. Such considerations disfavour HNLs with masses $m_{N}\lesssim 1$ GeV, for small active-sterile mixing strengths required by current constraints and expected for light neutrino mass generation. Even lighter and longer-lived $N$ are likewise disfavoured as they will act as additional degrees of freedom, or inject them through their decays, and may overclose the universe.
@@
The above considerations apply for the minimal scenario where the HNL only couples via the active-sterile mixing. Instead, HNLs may also interact with an expanded exotic sector within specific beyond-the-SM scenarios. Often, these are high-scale scenarios where the HNL is charged under an additional gauge force, such as $B-L$ [17, 18, 19, 20, 21] and left-right-symmetric [22, 23, 24, 25, 26, 19] models. Less explored is the possibility of coupling the HNL to a lighter exotic sector. We here consider such a scenario where the HNL couples to a light, exotic pseudoscalar $a$, akin to an axion-like particle (ALP) [27, 28, 29]. Recent works exploring similar HNL-ALP portals in colliders can be found in [30, 31, 32, 33] and [34, 32] in beam dump experiments. We will refer to our dark scalar as ALP in the following, and it is dark in the sense that it couples only to the HNL in the first instance, with interactions to the SM suppressed by the active-sterile mixing and through loops.
@@ 2.1 Lagrangian
$\mathcal{M}_{\nu}=\begin{pmatrix}0&\mathcal{M}_{D}\\ \mathcal{M}_{D}^{T}&\mathcal{M}_{R}\end{pmatrix},$ (2.2)
@@
$U=\begin{pmatrix}U_{\nu\nu}&U_{\nu N}\\ U_{N\nu}&U_{NN}\end{pmatrix},$ (2.3)
@@
$U^{\dagger}\mathcal{M}_{\nu}U^{\ast}=U^{\dagger}\cdot\begin{pmatrix}0&\mathcal{M}_{D}\\ \mathcal{M}_{D}^{T}&\mathcal{M}_{R}\end{pmatrix}\cdot U^{\ast}=\begin{pmatrix}(m_{\nu})_{3\times 3}&0\\ 0&(M_{N})_{\mathcal{N}\times\mathcal{N}}\end{pmatrix}.$ (2.4)
@@
In addition to the regular seesaw sector, we also assume that the ALPs $a$ couples with the HNL $N$, but not with other SM particles directly. Hence, due to the pseudoscalar nature of ALPs, the Lagrangian for the ALP-HNL interaction is given by the derivative coupling [35] in the mass basis as
$\mathcal{L}_{aNN}=\sum_{\kappa=1}^{\mathcal{N}}\frac{1}{f_{a}}(\partial_{\mu}a)\bar{N}_{\kappa}\gamma^{\mu}\gamma_{5}N_{\kappa}=-\sum_{\kappa=1}^{\mathcal{N}}\frac{2i}{f_{a}}m_{N_{\kappa}}a\bar{N}_{\kappa}\gamma_{5}N_{\kappa},$ (2.6)
where $f_{a}$ is the ALP decay constant and $m_{N_{\kappa}}$ are the eigenvalues of $M_{N}$. The second equality holds by applying the equation of motion for the HNLs and removing a total derivative after integration by parts [36]. Due to the small active-sterile mixing $U_{\nu N}=U_{N\nu}\equiv\sqrt{m_{\nu}M_{N}^{-1}}$, an interaction between the ALP and the active neutrinos $\nu_{\lambda}$ is induced. In the mass basis of the active neutrinos and HNLs, we can write the interaction Lagrangian between the ALP and active neutrinos as
@@
$\approx-\sum_{\lambda,\lambda^{\prime}=1}^{3}\frac{2i}{f_{a}}m_{N}a|U_{N\nu}U_{\nu\nu}^{\ast}|^{2}_{\lambda\lambda^{\prime}}\bar{\nu}_{\lambda}\gamma_{5}\nu_{\lambda^{\prime}}$ (2.7)
@@
$\mathcal{L}_{aN\nu}=-\sum_{\lambda^{\prime}=1}^{3}\sum_{\kappa=1}^{\mathcal{N}}\frac{2i}{f_{a}}m_{N_{\kappa}}a(U_{N\nu}^{\ast}U_{\nu\nu})_{\kappa\lambda^{\prime}}\bar{N}_{\kappa}\gamma_{5}\nu_{\lambda^{\prime}}.$ (2.8)
While at least two HNLs are needed to explain all active neutrino masses and mixing, we are mainly interested in elucidating the principle effects in our framework. For simplicity, by considering a single HNL, $N_{1}\equiv N$, we can acquire all the important information from the framework using this simplified viewpoint. Furthermore, we consider the active-sterile mixing of $N_{1}$ with a single active neutrino $\nu_{1}$ which is mostly electron-type neutrino i.e., $\nu_{1}\equiv\nu_{e}$, without loss of generality. From now on, we take $|U_{\nu\nu}|^{2}\sim\mathcal{O}(1)\gg|U_{\nu N}|^{2}=|U_{N\nu}|^{2}$ and $U_{\nu N}=U_{eN}$. Equations (2.6) and (2.7) then take the form
@@
$\mathcal{L}_{a\nu\nu}$ $=-\frac{2i}{f_{a}}m_{N}|U_{eN}|^{2}a\bar{\nu}_{e}\gamma_{5}\nu_{e}=-\frac{2i}{f_{a}}m_{\nu}a\bar{\nu}_{e}\gamma_{5}\nu_{e},$ (2.9)
@@ 2.2 HNL decays
A Majorana HNL with a mass of a few hundred MeV to GeV decays via various channels into SM particles. All these decays are four-fermion interactions mediated by either a $Z$ or $W$ boson. We will briefly discuss these SM decay channels of HNLs here with a detailed discussion found in [14, 15]. In our framework with only one HNL species that couples with the first generation only, we have the corresponding decay channels of HNLs as:
@@
$\Gamma^{\nu_{e}\ell^{-}\ell^{+}}=|U_{eN}|^{2}\frac{G_{F}^{2}m_{N}^{5}}{96\pi^{3}}\left[\left(C_{1}+2\sin^{2}\theta_{W}\delta_{e,\ell}\right)f_{1}(x_{\ell})+\left(C_{2}+\sin^{2}\theta_{W}\delta_{e,\ell}\right)f_{2}(x_{\ell})\right],$ (2.10)
@@
$C_{1}=\frac{1}{4}(1-4\sin^{2}\theta_{W}+8\sin^{4}\theta_{W}),\qquad C_{2}=\frac{1}{2}(-\sin^{2}\theta_{W}+2\sin^{4}\theta_{W}),$ (2.11)
@@
$f_{2}(x)$ $=4\left[x^{2}(2+10x^{2}-12x^{4})\sqrt{1-4x^{2}}+6x^{4}(1-2x^{2}+2x^{4})L(x)\right],$ (2.12)
@@
$L(x)=\ln\left[\frac{1-3x^{2}-(1-x^{2})\sqrt{1-4x^{2}}}{x^{2}(1+\sqrt{1-4x^{2}})}\right].$ (2.13)
@@
$\Gamma^{e^{-}\ell^{+}\nu_{\ell}}=|U_{eN}|^{2}\frac{G_{F}^{2}m_{N}^{5}}{192\pi^{3}}\left[1-8x_{\ell}^{2}+8x_{\ell}^{6}-x_{\ell}^{8}-12x_{\ell}^{4}\ln(x_{\ell}^{2})\right].$ (2.14)
@@
$\Gamma^{P\nu_{e}}=\frac{G_{F}^{2}m_{N}^{3}}{32\pi}f_{P}^{2}|U_{eN}|^{2}(1-x_{P}^{2})^{2},$ (2.16)
with the meson decay constant $f_{P}$ whose values we have taken from [15] and $x_{P}=m_{P}/m_{N}$.
@@
$\Gamma^{P^{+}e^{-}}=\frac{G_{F}^{2}m_{N}^{3}}{16\pi}f_{P}^{2}|U_{eN}|^{2}|V_{qq^{\prime}}|^{2}\lambda^{1/2}(1,x_{P}^{2},x_{e}^{2})\left[1-x_{P}^{2}-x_{e}^{2}(2+x_{P}^{2}-x_{e}^{2})\right].$ (2.17)
@@
$\Gamma^{V\nu_{e}}=\frac{G_{F}^{2}m_{N}^{3}}{32\pi m_{V}^{2}}f_{V}^{2}\kappa_{V}^{2}|U_{eN}|^{2}(1+2x_{V}^{2})(1-x_{V}^{2})^{2}.$ (2.18)
@@
$\Gamma^{V^{+}e^{-}}=\frac{G_{F}^{2}m_{N}^{3}}{16\pi m_{V^{\pm}}^{2}}f_{V}^{2}|U_{eN}|^{2}|V_{qq^{\prime}}|^{2}\lambda^{1/2}(1,x_{V}^{2},x_{e}^{2})\left[(1-x_{V}^{2})(1+2x_{V}^{2})+x_{e}^{2}(x_{V}^{2}+x_{e}^{2}-2)\right].$ (2.19)
@@
$+\sum_{P}\Gamma^{P\nu_{e}}+\sum_{P}2\Gamma^{P^{+}e^{-}}+\sum_{V}\Gamma^{V\nu_{e}}+\sum_{V}2\Gamma^{V^{+}e^{-}}.$ (2.20)
@@
Due to the interaction in Eq. (2.9), a new decay channel for HNLs is also possible, where the HNL decays as $N\to a\nu$ via the active-sterile mixing at tree level. The decay width, in the limit where $m_{\nu}\ll m_{N}$, can be written as
$\Gamma^{N\to a\nu}=\frac{|U_{eN}|^{2}m_{N}^{3}}{4\pi f_{a}^{2}}\sqrt{1+\left(\frac{m_{a}}{m_{N}}\right)^{2}}\left[1-\left(\frac{m_{a}}{m_{N}}\right)^{2}\right]^{3/2}\approx\frac{|U_{eN}|^{2}m_{N}^{3}}{4\pi f_{a}^{2}},$ (2.21)
@@
$\tau_{N\to a\nu}\approx 8.6\times 10^{-4}~{}\text{s}\times\left(\frac{f_{a}}{1~{}\text{TeV}}\right)^{2}\times\left(\frac{10^{-14}}{|U_{eN}|^{2}}\right)\times\left(\frac{1~{}\text{GeV}}{m_{N}}\right)^{3}.$ (2.22)
The total HNL decay width to SM particles can be approximately written as [9]
@@
$\approx\left(\frac{6f_{M}^{2}}{\pi}+\frac{m_{N}^{2}}{20\pi^{3}}\right)m_{N}^{3}|U_{eN}|^{2}G_{F}^{2},$ (2.23)
where $f_{M}\approx\mathcal{O}(0.1)$ GeV corresponds to the typical decay constant of the produced pseudoscalar or vector meson. Comparing this with Eq. (2.21), it is clear that the $N\to a\nu$ channel will be dominant if
$m_{N}<\frac{\sqrt{5|1-24G_{F}^{2}f_{M}^{2}f_{a}^{2}|}\pi}{f_{a}G_{F}},$ (2.24)
@@
The branching ratios of the HNL decaying to various particles as a function of mass of the HNL for two different values of $f_{a}$ are shown in Fig. 1. In this work, as we are mostly interested in the regime where $m_{a}\ll m_{N}$, correspondingly we have chosen $m_{a}=1$ keV and varied $m_{N}\in[10^{-3}-50]$ GeV range. In the left panel, we have considered $f_{a}=1$ TeV where the Br$(N\rightarrow a\nu)\approx 100\%$ (solid green) throughout the entire mass range of HNL, while the behaviour of different SM decay channels are shown as, $N\rightarrow\nu\nu\nu$ (solid black), $N\rightarrow\nu_{e}(e^{+}e^{-}+\mu^{+}\mu^{-}+\tau^{+}\tau^{-})$ (solid orange), $N\rightarrow e^{-}(\nu_{\mu}\mu^{+}+\nu_{\tau}\tau^{+})$ (solid magenta), different pseudoscalar ($P$) and vector ($V$) mesonic channels (which become relevant for $m_{N}\geq m_{\pi}$) as $N\rightarrow P\nu$ (solid red), $N\rightarrow Pe$ (dashed brown), $N\rightarrow V\nu$ (solid cyan), $N\rightarrow Ve$ (dashed blue). In the right panel of the figure we have taken $f_{a}=10^{2.5}$ TeV, for which the coupling of the ALP to HNL significantly decreases. Here, for lower mass of HNL the axionic decay channel still dominates the scenario, as expected from eq. (2.24), while for $m_{N}>m_{\pi}$, as the mesonic channels come into picture, the branching ratio to axionic channel drops significantly as compared to the SM decay channels.